# Jose Tupayachi

### v1

@@ 7 Conclusion & Recommendations
This study introduces TW-GCN, a specialized spatio-temporal framework designed to enhance the prediction of electric vehicle charging demand across Tennessee. By fusing temporal modeling (1DCNN, LSTM, GRU, and RNN) with spatial dependencies captured via GCNs, the architecture effectively learns both temporal sequences and regional interconnectivity. Through extensive experimentation with real-world EV charging data and multiple model configurationsincluding variations in lag hours, clustering strategies, and prediction horizons. We demonstrate the model’s capacity to balance short-term responsiveness with long-term stability. Key findings indicate that regional disparities in station availability and population density, especially in East Tennessee, can be meaningfully addressed by leveraging predictive performance at the ZIP or coordinates level. These insights equip decision makers with actionable intelligence for infrastructure deployment, policy alignment, and equitable EV accessibility.
Our integrated use of weather, traffic, and spatial POIs further enhances the realism and applicability of the forecasting outputs, leading to a more robust modeling pipeline. Importantly, the short, mid and long horizon analyses show that trade-offs between volatility and trend awareness must be carefully managed, with the mid horizon offering a particularly balanced outcome in terms of error performance and operational relevance.
Building upon the results presented. First, expanding the dataset to include additional states or regions with different EV adoption rates can validate the model’s generalizability. Exploration of transformer-based architectures or attention mechanisms could enhance the ability to capture long-range dependencies across time and space. From a policy and infrastructure standpoint, integrating cost-benefit analysis into the modeling pipeline would support more precise investment decisions. Finally, addressing data gaps particularly and enhancing the granularity of station-level features (e.g., charging speed, stall availability) will further elevate the practical utility of the proposed approach.
Appendix A Distance Functions
Algorithm 1 DTW Distance Function
1:Function DTWDistance(s: array[1..n], t: array[1..m])
2:DTW := array[0..n, 0..m]
3:for $i:=0$ to $n$ do
4:  for $j:=0$ to $m$ do
5:    DTW[i, j] := $\infty$
6:  end for
7:end for
8:DTW[0, 0] := 0
9:for $i:=1$ to $n$ do
10:  for $j:=1$ to $m$ do
11:    cost := $d(s[i],t[j])$
12:    DTW[i, j] := cost + minimum(DTW[i-1, j],  DTW[i, j-1],  DTW[i-1, j-1])
13:  end for
14:end for
15:return DTW[n, m]
Algorithm 2 Haversine Distance Function
1:lat1, lon1, lat2, lon2 in degrees
2:d – Haversine distance in kilometers
3:Convert lat1, lon1, lat2, lon2 from degrees to radians
[…]
9:d $\leftarrow R\cdot c$ return d
[…]

### v3

@@ 7 Conclusion & Recommendations
This study introduces TW-GCN, a specialized spatio-temporal framework designed to enhance the prediction of electric vehicle charging demand across Tennessee. By fusing temporal modeling (1DCNN, LSTM, GRU, and RNN) with spatial dependencies captured via GCNs, the architecture effectively learns both temporal sequences and regional interconnectivity. Through extensive experimentation with real-world EV charging data and multiple model configurations, including variations in lag hours, clustering strategies, and prediction horizons. We demonstrate the model’s capacity to balance short-term responsiveness with long-term stability. Key findings indicate that regional disparities in station availability and population density, especially in East Tennessee, can be meaningfully addressed by leveraging predictive performance at the ZIP or coordinates level. These insights equip decision makers with actionable intelligence for infrastructure deployment, policy alignment, and equitable EV accessibility.
Our integrated use of weather, traffic, and spatial POIs further enhances the realism and applicability of the forecasting outputs, leading to a more robust modeling pipeline. Importantly, the short, mid, and long horizon analyses show that trade-offs between volatility and trend awareness must be carefully managed, with the mid horizon offering a particularly balanced outcome in terms of error performance and operational relevance.
Building upon the results presented, future work could expand the dataset to include additional states or regions with different EV adoption rates can validate the model’s generalizability. Exploration of transformer-based architectures or attention mechanisms could enhance the ability to capture long-range dependencies across time and space. From a policy and infrastructure standpoint, integrating cost-benefit analysis into the modeling pipeline would support more precise investment decisions. Finally, addressing data gaps particularly and enhancing the granularity of station-level features (e.g., charging speed, stall availability) will further elevate the practical utility of the proposed approach.
We note that the statements and findings presented in this study are solely those of the authors and do not necessarily reflect the views or positions of ChargePoint, Inc.